a) A point on the outer edge of a disk with a diameter of 1.8 m that reaches 357 revolutions per minute. Calculate the angular acceleration assuming the disc started from rest and it took 21 seconds to achieve 357 rpm. Express your answer correct to 2 decimal places. Add the appropriate units.
b) How many rotations does this disc do in 21 seconds? Express your answer as a whole number and add the appropriate units (rotations).
1
Expert's answer
2018-04-27T14:36:07-0400
a) The angular acceleration is
α=ω/t=2π/60 357/21=1.78 rad/s^2 .
b)
N=ω t/2=357/60 21/2=62 revolutions.